OpenStudy (ranya99):

If x² + 8x +c is a perfect square trinomial, what is the value of c? a. 32 b. 8 c. 16 d. 4 _______________________________ Find the values for x in the following equation: x² + 4x -32 = 0 a. -8 & 4 b. 8 & -4 c. 8 & 4 d. 0, -8, & 4 ________________________________ Find the x-intercepts for the following equation: y = 6x² - 7x - 3 a. (3/2,0) & (-1/3,0) b. (3,0) & (-1,0) c. (7,0) & (12,0) d. (-3/2,0) & (1/3,0) ______________________________ Find the values of y for the following equation: 2y² - 5y + 2 = 5 a. y = 1/2 & -3 b. y = 1/2 & 2 c. y = -1/2 & -2 d. y = -1/2 & 3

OpenStudy (anonymous):

Okay @Ranya99, you're going to have to do some of the work here. We're here to help, not just do your homework for you. What are you stuck on? Answer these questions please: 1) What is a perfect square trinomial? 2) How do we find x intercepts of functions?

OpenStudy (ranya99):

yeah I don't just need answers, I need to understand them too. Both acually.

OpenStudy (anonymous):

Okay no problem, just in future try to say what it is you want rather than just post questions and you'll be able to get the best help. Can you answer my previous questions? If you don't know the answers, it's quite easy to look up :)

OpenStudy (ranya99):

a perfect square is a number that can be expressed as the product of two equal integers to find x-intercept, plug in 0 for y and solve for x

OpenStudy (anonymous):

Exactly, so for 1 we want to find an a such that \[(x+a)^2=x^2+8x+c\]It would normally be something of the form (ax+b)^2 but we know the a in that case would be 1 so we can ignore it. \[(a+x)^2=x^2+2ax+a^2=x^2+8x+c\] We know 2a=8 and a^2=c. What are a and c?

OpenStudy (anonymous):

Meant to write (x+a)^2 not (a+x)^2 - I know they're the same, it's just confusing.

OpenStudy (ranya99):

I'm confused

OpenStudy (anonymous):

Okay so as you said, a perfect square trinomial is the square of two products. (ax+b)^2. When we expand this, we get\[a^2x^2+2abx+b^2\]Setting this equal to your equation gives us\[a^2x^2+2abx+b^2=x^2+8x+c\]Now we equate coefficients to solve this. In front of the x^2, we have a^2 on the left and 1 on the right. So a^2=1 therefore a=1 (this is the step I skipped before). In front of x, we have 2ab on the left and 8 on the right. We know a=1 from before, so our equation is 2b=8, so b=4. Now we're left with b^2=c, and we just found b=4 so c=4^2=16. This gives us: \[(x+4)^2=x^2+8x+16\] Does that explain it?

OpenStudy (ranya99):

yeah

OpenStudy (anonymous):

The second question doesn't really make sense to me, is that exactly how you see it?

OpenStudy (ranya99):

yes

OpenStudy (anonymous):

Ill call a haxor to help out with this.

OpenStudy (anonymous):

@FlappyFlapper

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yes, I guess

OpenStudy (anonymous):

Yes I guess, what a Mess

OpenStudy (anonymous):

We will both say what at the same time: 3, 2, 1

OpenStudy (anonymous):

Oh sorry, I'm being thick. How do you solve:\[x^2 + 4x -32 = 0 \]

OpenStudy (anonymous):

What

OpenStudy (anonymous):

Yes I guess, a dress that's a mess.

OpenStudy (anonymous):

With a bess to obsess

OpenStudy (ranya99):

@tom982 I solved this one, it's -8 & 4

OpenStudy (anonymous):

Correct. Now for 3 it's the same, just set y=0 and solve for x. For 4, we rearrange to make it equal 0 then solve like we did in 2 and 3.

OpenStudy (ranya99):

@tom982 alright, got it thanks a lot

OpenStudy (anonymous):

No problem, happy to help.